The head on a spillway is 1.5m and the discharge is 100m3 /min. What is the head on the scale model of 1:80? Determine the model discharge. 


1
Expert's answer
2021-08-16T02:19:22-0400

Head

"\frac{h_P}{h_m}=L_r\\
\frac{1.5}{h_m}=80\\
h_m= \frac{1.5}{80}\\
h_m= 0.01875 m"


Discharge

"\frac{Q_p}{Q_m}=(Lr)^{2.5}\\
\frac{100}{Q_m}=(80)^{2.5}\\
Q_m= \frac{100}{80^{2.5}}\\
Q_m= 1.747* 10^{-3}m ^3/mm"